# Multi-Timescale Gradient Sliding for Distributed Optimization

### v1

@@ 1.2 Contributions
We make the following contributions to the convex non-smooth distributed optimization literature.
Formulation: generalized, block-decomposable penalties for consensus constraints.
We propose relaxing the consensus constraints through generic convex penalty functions (($\mathcal{P}_{r}$)), generalizing the characteristic function penalty used in previous works. Motivated by settings where the consensus constraints consist of orthogonal components (e.g. the hierarchical setting), we propose using block-decomposable penalties. This results in a saddle point formulation ($\mathcal{P}_{s}$) with decomposable duals, a structure which we exploit in designing the multi-timescale algorithms.
Further, we propose a set of conditions on the growth rates of penalties (Lemma 2.3 and Corollary 2.1), under which the duality gap for ($\mathcal{P}_{s}$) provides upper bounds on objective value suboptimality and consensus constraint violation. This relates solution qualities for the penalized problems (($\mathcal{P}_{r}$) and ($\mathcal{P}_{s}$)) back to the original distributed optimization problem ($\mathcal{P}$).
Algorithm: (accelerated) multi-timescale gradient sliding.
We propose MT-GS and AMT-GS for saddle point problems with block-decomposable duals. At each iteration, our algorithms apply a generalized communication sliding procedure [lan_communication-efficient_2020] – consisting of multiple subgradient (mirror descent) steps – to approximate the proximal updates of the primal variables. By comparison, the dual blocks are updated at potentially different, user-chosen frequencies. Denoting the (weighted) average of these frequencies as $\overline{r}$, we show that the duality gap of MT-GS converges at the rate of $O(\overline{r}/k)$ for Lipschitz convex objectives (Lemma 3.1, Theorem 3.1), and the duality gap of AMT-GS converges at the rate of $O((\overline{r})^{2}/\mu k^{2})$ for $\mu$-strongly convex objectives (Lemma 4.1, Theorem 4.1). When specialized in the setting of distributed optimization, to find an $\epsilon$-suboptimal solution, the communication round and subgradient oracle complexities of our algorithms (Table 1) have optimal dependency on $\epsilon$.
condition (1) communication round subgradient oracle
MT-GS (Corollary 3.2, 3.3) $\mu\geq 0$ $O(\frac{\overline{r}A\sqrt{D^{X}}}{\epsilon})$ $O(\frac{\overline{r}mM^{2}D^{X}}{\epsilon^{2}})$
AMT-GS (Corollary 4.2, 4.3) $\mu>0$ $O(\frac{\overline{r}A}{\sqrt{\mu\epsilon}})$ $O(\frac{\overline{r}\sqrt{m}M^{2}}{\mu\epsilon})$
Table 1: Communication rounds and subgradient oracles needed to find an $\epsilon$ suboptimal solution to ($\mathcal{P}$). $\overline{r}=\sum_{s=1}^{S}r_{s}\rho_{s}$, $A$ measures function similarities, $D^{X}=D_{w^{X}}(X^{*},X^{init})$. Subgradient oracle for AMT-GS assumes $S=1$; for general $S$, see Section 4.2.
To the best of our knowledge, MT-GS and AMT-GS are the first deterministic algorithms for saddle point problems with block-decomposable duals that allow different number of updates for different dual blocks. This extra flexibility in choosing the update frequencies allows one to design more communication-efficient algorithms, especially when the cost of updating different duals and/or the domain sizes of different duals are different (Section 3.5).
Beyond Lipschitz constants: function similarity based complexity.
We formalize the notion of function similarity for non-smooth objectives (Definition 2.1) studied in [ArjevaniShamir2015Communication], and show that with proper choices of the penalties, the communication round complexities of MT-GS and AMT-GS have linear, and thus optimal, dependency on function similarities (Sections 3.5 and 4.2). This provides positive answers to the open question whether the theoretical communication round lower bounds proposed in [ArjevaniShamir2015Communication] can be attained.
@@ 1.3 Roadmap
In Section 2, we introduce the penalty and primal-dual formulation (($\mathcal{P}_{r}$) and ($\mathcal{P}_{s}$)) and propose a set of requirements on the penalty functions (Lemmas 2.2 and 2.3, and Corollary 2.1). We also formalize the notion of function similarities for non-smooth objectives (Definition 2.1). In Section 3, we present our MT-GS algorithm and provide its convergence guarantees for general saddle point problems (Section 3.3) and for ($\mathcal{P}$) (Section 3.4). The section ends with further discussion on the cost and the dependence on function similarities (Section 3.5). In Section 4, we give the convergence results of our AMT-GS for general saddle point problems. Then we propose a communication-free initialization process (Section 4.1), and discuss the overall communication round complexities (Section 4.2). In Section 5, we present numerical experiments for support vector machine problems with regularized hinge losses. Finally, in Section 6, we point out directions of future work.
2 Setup and formulations

### v2

@@ 1.2 Contributions
We propose multi-timescale PDHG for saddle point problems with block-decomposable duals, where different dual blocks are updated at different rates.
  • To ensure convergence with arbitrary updating rates, we propose novel multi-timescale extrapolation steps for the dual updates (5) and mixtures of Bregman divergence for the primal updates (6). The duality gaps of our algorithms converge at the optimal rates: $O(1/k)$ and $O(1/k^{2})$ for general and strongly convex objectives, respectively, despite the potentially outdated information in the dual.
  • To the best of our knowledge, for the saddle point problems considered, our multi-timescale PDHG algorithms are the first deterministic updating mechanisms which still allow different (arbitrary) updating rates for different dual blocks.
  • We quantify how updating rates influence convergence: the duality gaps show linear dependencies on weighted averages of the updating rates (Corollary 3.3) and its square (Corollary 3.5) for general and strongly convex objectives, respectively. Thus, our algorithms are robust to the maximum updating rates. [^2] We call the inverses of frequencies the rates. Thus, the largest updating rate corresponds to the smallest frequency.
We apply our algorithms to convex non-smooth distributed optimization. More specifically, we propose relaxing the consensus constraints via generic convex, block-decomposable penalty functions (Lemma 4.3 and Corollary 4.1), which generalize the characteristic-function penalty used in prior work. This yields saddle-point formulations ($\mathcal{P}_{d}^{lift}$) with block-decomposable dual variables, enabling our multi-timescale PDHG method, which is particularly well suited to settings with heterogeneous communication costs among agents.
Moreover, we show that with proper choices of the penalties, the communication round complexities of our algorithms have linear, and thus optimal, dependency on similarities between gradients of the local objectives. This provides positive answers to the open question whether the theoretical communication round lower bounds proposed in [4] can be attained.
@@ 1.3 Roadmap
In Section 2, we present additional details and assumptions about the saddle point problem ($\mathcal{P}_{b}$). Then, in Section 3.1, we propose our multi-timescale PDHG for ($\mathcal{P}_{b}$), and in Sections 3.2 and 3.3, we present the convergence properties of our algorithms with and without strong convexity, respectively.
In Section 4, we apply our multi-timescale PDHG to distributed optimization problems: we first show that ($\mathcal{P}_{d}$) can be reformulated in the lifted space in the form of ($\mathcal{P}_{b}$) (Section 4.1) and propose a set of conditions on the consensus constraints (Section 4.2) and penalties (Section 4.3). The requirements on the penalties motivate the definition of function similarities (Definition 4.1). Then, we describe how multi-timescale PDHG can be applied and the communication involved (Section 4.4), and provide the convergence results in Sections 4.5 and 4.6.
Numerical experiments are provided in Section 5.
2 Setup
For the problem ($\mathcal{P}_{b}$), we assume $\mathbb{R}^{d}$ is equipped with a norm $\|\cdot\|$ not necessarily generated by the Euclidean inner product, and we equip $\mathcal{X}$ with a distance generating function [^3] For a convex closed set $\mathcal{S}$, a function $w:\mathcal{S}\to\mathbb{R}$ is a distance generating function [20] with modulus $\nu>0$ w.r.t. $\|\cdot\|$ if $w$ is continuously differentiable and $\langle x-z,\nabla w(x)-\nabla w(z)\rangle\geq\nu\|x-z\|^{2},\quad\forall x,z\in\mathcal{S}.$ The Bregman divergence generated by $w$ is defined as $D_{w}(x,z):=w(x)-w(z)-\langle\nabla w(z),x-z\rangle$. $w_{X}:\mathcal{X}\to\mathbb{R}$ with modulus $1$.
We further denote $\mathcal{Y}_{s}:=\mathsf{dom}(R_{s}^{*})=\{y_{s}\in\mathbb{R}^{n_{s}},~R_{s}^{*}(y_{s})<\infty\}$ as the domain of $R_{s}^{*}$, and define $R:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ as $R(Y)=\sum_{s=1}^{S}R_{s}(y_{s})$. It’s easy to see that $R^{*}(Y)=\sum_{s=1}^{S}R_{s}^{*}(y_{s})$ and $\mathcal{Y}:=\mathsf{dom}(R^{*})=\prod_{s=1}^{S}\mathsf{dom}(R_{s}^{*})$. Similarly, for each $s\in[S]$, we assume that $\mathbb{R}^{n_{s}}$ is equipped with a norm $\|\cdot\|$ not necessarily generated by the Euclidean inner product, and we equip $\mathsf{dom}(R_{s}^{*})$ with distance generating function $w_{y_{s}}:\mathsf{dom}(R^{*}_{s})\to\mathbb{R}$ with modulus $1$, and $w_{Y}(Y):=\sum_{s=1}^{S}w_{y_{s}}(y_{s})$.
For convenience, we define $K:\mathbb{R}^{d}\to\mathbb{R}^{n}$ as $(KX)_{s}=K_{s}X$ for each $s\in[S]$. Thus, ($\mathcal{P}_{b}$) can be compactly written as
$\min_{X\in\mathcal{X}}\max_{Y\in\mathbb{R}^{n}}F(X)+\langle KX,Y\rangle-R^{*}(Y).$ ($\mathcal{P}_{c}$)
In addition, since $R_{s}$ is convex and lower-semicontinuous, $R_{s}=R_{s}^{**}$ (Theorem 11.1 [39]), and importantly,
$R_{s}(K_{s}X)=\sup_{y_{s}\in\mathbb{R}^{n_{s}}}\langle K_{s}X,y_{s}\rangle-R_{s}^{*}(y_{s}).$
Thus, ($\mathcal{P}_{b}$) is also equivalent to the following primal-only formulation:
$\min_{X\in{\mathcal{X}}}F(X)+\sum_{s=1}^{S}R_{s}(K_{s}X).$ ($\mathcal{P}_{p}$)
In the rest of the work, when there is no confusion on the domain of the function, we abbreviate $w_{X},w_{Y},w_{y_{s}}$ as $w$ and the associated Bregman divergence as $D$. We also make the following assumption.
Assumption 2.1.
$\min_{y_{s}\in\mathsf{dom}(R^{*}_{s})}R_{s}^{*}(y_{s})+\langle g,y_{s}\rangle+D(y_{s},\overline{y}_{s}).$
For any $g\in\mathbb{R}^{d}$, the following problem can be solved exactly:
$\min_{X\in\mathcal{X}}\langle g,X\rangle+w(X).$
Performance measure. We use the duality gap to measure the performance of $Z=(X,Y)$. More precisely, we define $G:\mathcal{Z}\times\mathcal{Z}\to\overline{\mathbb{R}}$ where $\mathcal{Z}=\mathcal{X}\times\mathbb{R}^{n}$ as
$G(X,Y;X^{\prime},Y^{\prime}):=\langle KX,Y^{\prime}\rangle+F(X)-R^{*}(Y^{\prime})-\left\{\langle KX^{\prime},Y\rangle+F(X^{\prime})-R^{*}(Y)\right\},$ (2)
and we will provide upper bounds on $G(Z):=\sup_{Z^{\prime}\in\mathcal{Z}}G(Z;Z^{\prime})$ where $Z$ is the output of our algorithms.
3 Multi-timescale primal-dual updates
To solve the saddle point problem ($\mathcal{P}_{c}$), one of the most popular algorithms is the Primal-Dual Hybrid Gradient (PDHG) algorithm[14, 27, 13], which updates the primal variable $X$ and the dual variable $Y$ iteratively. At iteration $k=0,1,\ldots$,
$\widetilde{X}^{k}$ $=X^{k-1}+\alpha_{k}(X^{k-1}-X^{k-2})$ (3a)
$Y^{k}$ $=\operatorname*{argmin}_{Y\in\mathbb{R}^{n}}R^{*}(Y)+\langle-K\widetilde{X}^{k},Y\rangle+\tau_{k}D(Y,Y^{k-1})$ (3b)
$X^{k}$ $=\operatorname*{argmin}_{X\in\mathcal{X}}F(X)+\langle K^{*}Y^{k},X\rangle+\eta_{k}D(X,X^{k-1}),$ (3c)
where $\tau_{k},\eta_{k}>0$ are parameters depending on the operator norm of $K$, and $\alpha_{k}\in[0,1]$. The convergence rates of PDHG under various assumptions of the objective functions $F$ and $R$ have been well established [14]. For instance, for general convex $F$ and $R$, the duality gap (of the ergodic mean) converges at the rate of $O(1/k)$; further, if $F$ is strongly convex, the accelerated rate of $O(1/k^{2})$ can be achieved.
Approximation to (3c) through gradient sliding and auxiliary primal sequence. For a generic (convex, potentially non-linear) objective function $F$, it’s sometimes unreasonable to assume that one can find the exact minimizer in the update (3c). Indeed, in Assumption 2.1, we only assume that $\langle g,X\rangle+w(X)$ can be exactly minimized. Thus, [28, 27] propose and use gradient sliding techniques, which approximate the minimizer through multiple iterations of mirror descent. In addition, to ensure the convergence of the overall PDHG algorithm with inexact updates, an auxiliary sequence $\widehat{X}^{k}$ is constructed, and the right hand side of (3a) is replaced with $X^{k-1}+\alpha_{k}(\widehat{X}^{k-1}-X^{k-2})$.
Block-decomposable dual updates. In ($\mathcal{P}_{b}$), both $R^{*}(Y)=\sum_{s=1}^{S}R^{*}_{s}(y_{s})$ and $D(Y,Y^{k-1})=\sum_{s=1}^{S}D(y_{s},y_{s}^{k-1})$ are block-decomposable. Thus, the dual update (3b) is also block-decomposable:
$y_{s}^{k}=\operatorname*{argmin}_{y\in\mathbb{R}^{n_{s}}}R_{s}^{*}(y_{s})+\langle-K_{s}\widetilde{X}^{k},y_{s}\rangle+\tau_{k}D(y_{s},y_{s}^{k-1}),\quad s=1,\ldots,S.$ (4)
This makes block-coordinate descent type of algorithms possible. As an example, S-PDHG [13] updates a random subset of the dual blocks at each iteration. The flexibility in choosing the sampling distribution allows one to control the frequency of updates of different blocks. However, due to the randomness, the $O(1/k)$ rate of convergence is shown only for the expected duality gap.
To maintain the deterministic convergence guarantee as well as the flexibility in choosing the number of updates applied to each dual block, we propose a multi-timescale updating mechanism for ($\mathcal{P}_{b}$), where different dual blocks are updated at potentially different rates. More precisely, denoting the global time using $k=0,1,\ldots$, then the dual block $y_{s}$ is updated only at iteration $k=0,r_{s},2r_{s},\ldots$ for some positive integer $r_{s}$ and remains fixed for all other iterations. Due to this multi-timescale mechanism, two challenges arise.
  1. 1.
    Information delay in $y_{s}^{k}$. In the primal update (3a), the term $K^{*}Y^{k}=\sum_{s=1}^{S}K^{*}_{s}y_{s}^{k}$ depends on $y_{s}^{k}$, yet $y_{s}^{k}=y_{s}^{\lfloor k/r_{s}\rfloor\times r_{s}}$ contains information only till iteration $\lfloor k/r_{s}\rfloor\times r_{s}\leq k$, which could be “outdated”. To mitigate the negative effect of “information delay”, we propose using mixtures of Bregman divergences in (3c) to control how fast the primal sequence varies ((6)).
  2. 2.
    Multi-timescale information aggregation of $X^{k}$. The dual updates (3b) and (4) are defined for each global time $k$, and $\widetilde{X}^{k}$ depends on $X^{k-2}$ and $X^{k-1}$ only. In the multi-timescale setting, one can expect that $\widetilde{X}^{k}$ should depend on primal sequences over longer intervals (the length of which could depend on $r_{s}$), and could be different for different dual blocks. We propose multi-timescale extrapolation which aggregates $X^{(k-2r_{s}):(k-1)}$ for block $s$ ((5)).
In the rest of this section, we state the exact updating procedure in Section 3.1, and provide convergence results in Sections 3.2 and 3.3 for general convex and strongly convex $F$ respectively.
3.1 Multi-timescale updating procedure
In our multi-timescale updating procedure, we use $k=0,1,\ldots,N$ to denote the global time. Motivated by gradient sliding procedures in [27], we also allow approximate minimizers to the primal updates, and keep track of the pair $(X^{k},\widehat{X}^{k})$, which is assumed to satisfy the condition (7). This can be achieved using a generalized gradient sliding technique (Appendix A).
Initialization. We assume access to initializations $X^{init}\in\mathcal{X}$ and $y_{s}^{init}\in\mathcal{Y}_{s}$ for all $s\in[S]$, and we initialize $X^{k^{\prime}}=\widehat{X}^{k^{\prime}}=X^{init}$ and $y_{s}^{k^{\prime}}=y_{s}^{init}$ for all $k^{\prime}<0$.
Dual updates. We associate each dual $y_{s}$ with a rate $r_{s}\in\mathbb{N}$ and a local time $i_{s}=0,1,\ldots,N_{s}-1$, such that $N+1=r_{s}N_{s}$. For $k=0,1,\ldots,N$, $y_{s}$ remains dormant unless $k=r_{s}i_{s}$ for some $i_{s}\in\{0,1,\ldots,N_{s}-1\}$, where $y^{i_{s}}_{s}$ is computed as follows (which replaces (3a) and (3b)):
$\widetilde{X}_{s}^{i_{s}}$ $=\alpha_{s,i_{s}}\left(\sum_{k^{\prime}=r_{s}i_{s}-r_{s}}^{r_{s}i_{s}-1}\theta_{k^{\prime}}(\widehat{X}^{k^{\prime}}-X^{k^{\prime}-r_{s}})\right)+\sum_{k^{\prime}=r_{s}i_{s}-r_{s}}^{r_{s}i_{s}-1}\theta_{k^{\prime}+r_{s}}X^{k^{\prime}},$ (5a)
$y_{s}^{i_{s}}$ $=\operatorname*{argmin}_{y_{s}\in\mathbb{R}^{n_{s}}}\langle-\frac{1}{\sum_{k^{\prime}=r_{s}i_{s}}^{r_{s}i_{s}+r_{s}-1}\theta_{k^{\prime}}}K_{s}\widetilde{X}_{s}^{i_{s}},y_{s}\rangle+R^{*}_{s}(y_{s})+\tau_{s,i_{s}}D(y_{s},y^{i_{s}-1}_{s}).$ (5b)
Further, we denote $\overline{y}_{s}^{k}=y_{s}^{\lfloor k/r_{s}\rfloor}$, i.e. value of the dual block $y_{s}$ at the global time $k$, and we abbreviate $\overline{Y}^{k}=(\overline{y}_{s}^{k})_{s\in[S]}$.
Primal updates. First, for each $k$, we define
$\Phi^{k}(X):=F(X)+\langle K^{*}\overline{Y}^{k},X\rangle+\sum_{s=1}^{S}\eta_{k,s}D(X,X^{k-r_{s}}).$ (6)
All primal variables are updated at each global time as approximate minimizer to (6), such that there exists some $\delta_{k}:\mathcal{X}\to\mathbb{R}$ which could depend on $X^{k},X^{k-1}$ and other algorithm parameters, the following is satisfied:
$\Phi^{k}(\widehat{X}^{k})\leq\Phi^{k}(X)-(\frac{\mu}{C}+\sum_{s=1}^{S}\eta_{k,s})D(X,X^{k})+\delta_{k}(X),\quad\forall X\in\mathcal{X},$ (7)
where we assume (1) holds for some $\mu\geq 0$, and $D(X,X^{\prime})\leq\frac{C}{2}\|X-X^{\prime}\|^{2}$ for some $0<C\leq\infty$.
In case $\Phi^{k}$ can be minimized exactly, say $F(X)=\langle g_{F},X\rangle$ for some $g_{F}\in\mathbb{R}^{d}$ is a linear function, we can take $\widehat{X}^{k}=X^{k}=\operatorname*{argmin}_{X\in\mathcal{X}}\Phi^{k}(X)$ to be the exact minimizer, and then we can take $\delta_{k}(X)=0$ (Lemma A.1). Nevertheless, in (7), we also allow inexact minimizer. For generic convex objectives, this can be found through a generalization of the gradient sliding technique in [27]: we extend this technique from $S=1$ to $S\geq 1$ (see Appendix A for more details). More concretely, following Corollary A.1, the following results hold.
Corollary 3.1.
Consider the following updates using $GS$, the generalized gradient sliding procedure in Algorithm 3, where $T_{k}\in\mathbb{N}$ and $\eta_{k}=\sum_{s=1}^{S}\eta_{k,s}>0$, where $\eta_{k,s}\geq 0$,
$(X^{k},\widehat{X}^{k})=GS(F,{\mathcal{X}},D,T_{k},(\eta_{k,s})_{s\in[S]},(X^{k-r_{s}})_{s\in[S]},K^{*}\overline{Y}^{k},X^{k-1}).$ (8)
Assume that (1) holds with some $\mu\geq 0$, then with $\lambda_{t}=t+1$ and $\beta_{t}=\frac{t}{2}$ for $t\geq 1$, (7) holds with
$\delta_{k}(X)=\frac{2\eta_{k}}{T_{k}(T_{k}+3)}\left(D(X,X^{k-1})-D(X,X^{k})\right)+\frac{4M^{2}}{\eta_{k}(T_{k}+3)}.$
Further, if (1) holds with some $\mu>0$, and $D(X,X^{\prime})\leq\frac{C}{2}\|X-X^{\prime}\|^{2}$ for some $C<\infty$, then with $\lambda_{t}=t$ and $\beta_{t}=\frac{(t+1)\mu}{2\eta_{k}C}+\frac{t-1}{2}$, (7) holds with
$\delta_{k}(X)=\frac{2M^{2}/\eta_{k}}{T_{k}(T_{k}+1)}\sum_{t=1}^{T}\frac{\lambda_{t}}{\beta_{t}}.$
$(X^{k},\widehat{X}^{k}),~k=0,\ldots,17$$y_{1}^{i_{1}},~r_{1}=3,~i_{1}=0,\ldots,5$$y_{2}^{i_{1}},~r_{2}=3,~i_{2}=0,\ldots,5$$y_{3}^{i_{3}},~r_{3}=6,~i_{3}=0,\ldots,2$$y_{1}^{0}$$y_{1}^{1}$$y_{1}^{2}$$y_{1}^{3}$$y_{1}^{4}$$y_{1}^{5}$$y_{2}^{0}$$y_{2}^{1}$$y_{2}^{2}$$y_{2}^{3}$$y_{2}^{4}$$y_{2}^{5}$$y_{3}^{0}$$y_{3}^{1}$$y_{3}^{2}$
Figure 1: Updates for $S=3$, $r_{1}=r_{2}=3$ and $r_{3}=6$. Each marker represents one update: $(X^{k},\widehat{X}^{k})$ is updated at each global time $k=0,1,\ldots,17$. If generalized gradient sliding is used, then this involves $T_{k}$ iterations of mirror descent updates at iteration $k$. $y_{1}$ is updated at each local time $i_{1}=0,\ldots,5$, i.e. global time $k=0,3,6,\ldots,15$, and similarly for $y_{2}$ and $y_{3}$.
Algorithm 1 (Accelerated) Multi-timescale PDHG
$\{\alpha_{s,i_{s}}\},\{\theta_{k}\},\{\eta_{k,s}\},\{\tau_{s,i_{s}}\},\{r_{s}\},X^{init},Y^{init}$
Primal dual pair $Z^{N}$
Initialize $(X^{k^{\prime}},\widehat{X}^{k^{\prime}},Y^{k^{\prime}})\leftarrow(X^{init},X^{init},Y^{init})$ for all $k^{\prime}<0$